what is the answer to this math question?
x^2
___ -3+x= -3
7

Answer by josgarithmetic(39617)

Easier if you add +3 to both sides and then multiply both sides by 7. You should then have x^2+7x=0, much easier to solve.
x^2+7x = x(x+7) = 0
Either x=0 OR x=-7